OpenStudy (anonymous):

Find the domain of the function: f(x)= sqrt(x+8) / (x-3 Answer: (-inf, -8] U [3, inf) Is that the right answer folks?

OpenStudy (anonymous):

\[f(x)= \frac{ \sqrt{x+8}\ }{ x-3 }\]

OpenStudy (campbell_st):

well what value can't you have in the denominator or what value makes x - 3 = 0....? this will be the 1st restriction on the domain

OpenStudy (campbell_st):

next have a look at the numerator, I'll assume you are working in real numbers.. you can't have a negative square root... so you need to solve x + 8>=0

OpenStudy (campbell_st):

now you can start to decide what values can be in the domain

OpenStudy (anonymous):

still getting -8<=x<3 not sure i follow you campbell

OpenStudy (anonymous):

im trying to figure this out!

OpenStudy (anonymous):

@campbell_st so I'm getting a different answer: \[x \neq3, x \ge-8\] So my interval notation would be: [-8,3) U (3,infty)

OpenStudy (campbell_st):

thats the correct solution the denominator can be any real value except x = 3 and the numerator needs values greater than or equal to -8

OpenStudy (anonymous):

COOL!!! and there is the U (or) sign in the middle...

OpenStudy (campbell_st):

well thats set theory notation.... but makes sense,

OpenStudy (anonymous):

it says to enter the answer as either interval or union of intervals